# CYCle: Choosing Your Collaborators Wisely to Enhance Collaborative Fairness in Decentralized Learning

### v1

@@ Abstract
Collaborative learning (CL) enables multiple participants to jointly train machine learning (ML) models on decentralized data sources without raw data sharing. While the primary goal of CL is to maximize the expected accuracy gain for each participant, it is also important to ensure that the gains are fairly distributed. Specifically, no client should be negatively impacted by the collaboration, and the individual gains must ideally be commensurate with the contributions. Most existing CL algorithms require central coordination and focus on the gain maximization objective while ignoring collaborative fairness. In this work, we first show that the existing measure of collaborative fairness based on the correlation between accuracy values without and with collaboration has drawbacks because it does not account for negative collaboration gain. We argue that maximizing mean collaboration gain (MCG) while simultaneously minimizing the collaboration gain spread (CGS) is a fairer alternative. Next, we propose the CYCle protocol that enables individual participants in a private decentralized learning (PDL) framework to achieve this objective through a novel reputation scoring method based on gradient alignment between the local cross-entropy and distillation losses. Experiments on the CIFAR-10, CIFAR-100, and Fed-ISIC2019 datasets empirically demonstrate the effectiveness of the CYCle protocol to ensure positive and fair collaboration gain for all participants, even in cases where the data distributions of participants are highly skewed. For the simple mean estimation problem with two participants, we also theoretically show that CYCle performs better than standard FedAvg, especially when there is large statistical heterogeneity.
@@ 1 Introduction
Collaborative learning (CL) refers to a framework where several entities can work together by pooling their resources to achieve a common machine learning (ML) objective without sharing raw data. This approach offers particular advantages in domains like healthcare and finance that require access to extensive datasets, which can be difficult to acquire for any single entity due to the costs involved or privacy regulations such as HIPAA [3] and GDPR [10]. Federated learning (FL) [19] is a specific case of CL that enables multiple entities to collectively train a shared model by sharing their respective model parameters or gradients with a central server for aggregation. However, FL methods can lead to information leakage associated with sharing gradients/parameter updates with an untrusted third-party server that can potentially carry out reconstruction attacks [32, 30]. Furthermore, most CL algorithms assume that all participating agents are contributing equally to the learning process. When this assumption is violated, there is little incentive for collaboration because the gains are not distributed fairly. To avoid these pitfalls, a fully decentralized learning algorithm that fairly rewards the contributions of individual participants is required.
Recent works have attempted to tackle the challenge of collaborative fairness (CF) in FL settings. Shapley value (SV) [22] is often used to estimate marginal utility of participants in FL. However, computing the SV involves high computation and communication costs. To mitigate this problem, variants of SV have been proposed in [26, 16, 28]. But these methods only focus on FL with a central server. Existing private decentralized learning (PDL) algorithms such as CaPriDe [25], CaPC [7], and Cronus [4] emphasize on confidentiality, privacy and utility, while ignoring collaborative fairness.
In this work, we aim to bridge this significant gap by designing the CYCle protocol for knowledge sharing in the PDL setting. We focus on three goals - maximizing the mean accuracy gain across all participants, ensuring that no participant suffers performance degradation due to collaboration, and the accuracy gains are evenly distributed. To achieve these goals, we make the following contributions:
  • We analyze the collaborative fairness metric based on correlation coefficient between the contributions of participants and their respective final accuracies, and show that it fails in cases where the collaboration gain is negative.
@@ 2 Related Work and Background
Several challenges associated with CL have been addressed recently, including privacy [8, 7, 4, 25, 14, 9, 20], confidentiality [25, 7, 4], communication efficiency [19], robustness [17, 1, 2], and fairness [28, 31, 24, 18]. Most existing works focus only on FL algorithms with central orchestration such as FedAvg [19]. Due to the potential harm caused by gradient leakage in FL [30, 32], which can result in the disclosure of sensitive user data, there is a growing interest in decentralized learning algorithms that do not require centralized orchestration. Focus has been mainly on ensuring confidentiality, privacy, and utility in decentralized learning [25, 7, 4].
To achieve collaborative fairness in CL, it is critical to evaluate marginal contributions of participants [12, 28, 23, 13]. The main reason to have a robust data valuation technique is the heterogeneous nature of data distribution across clients, meaning that some clients may contribute more useful data than others. In vanilla FL, all participants receive the same global model irrespective of their contributions, which can potentially discourage collaboration and increase resistance to practical adoption. While Shapley Value (SV) can be used for data valuation, SV computation is expensive and hard to employ [22]. In CGSV [28], cosine similarity between local and global parameter updates is used to approximate the SV. The same cosine similarity approach is used in RFFL [27] for reputation scoring. Kumar et al. [16] used logistic regression models as proxies for client data and utilized an ensemble of them to approximate the SVs.
@@ 2.1 Preliminaries
Decentralized Learning: Let $N$ be the number of collaborating participants and $\mathcal{N}=\{1,2,...,N\}$. We assume that each participant $\mathcal{P}_{n},n\in\mathcal{N}$ has its own local training dataset $\mathcal{D}_{n}$ and a hold-out validation set $\tilde{\mathcal{D}}_{n}$. Based on the local training set, each participant can learn a local ML model $\mathcal{M}_{\theta_{n}}$ by mimimizing the local cross entropy loss $\mathcal{L}_{CE_{n}}$. The validation accuracy of this standalone local model $\mathcal{M}_{\theta_{n}}$ on $\tilde{\mathcal{D}}_{n}$ is denoted as $\mathcal{B}_{n}$. The goal in a typical FL algorithm is to build a single global model $\mathcal{M}_{\theta_{*}}$ that has a better accuracy than the standalone accuracy of all the local models. In contrast, the goal of each participant in decentralized learning is to obtain a better local model $\mathcal{M}_{\theta_{n}^{*}}$, which has a higher local validation accuracy $\mathcal{A}_{n}$ after collaboration. In other words, each participant $\mathcal{P}_{n}$ aims to maximize its collaborative gain (CG), which is defined as $\mathcal{G}_{n}=(\mathcal{A}_{n}-\mathcal{B}_{n})$. This is typically achieved by minimizing the following objective:
$\mathcal{L}_{n}=\mathcal{L}_{CE_{n}}+\lambda_{n}\mathcal{L}_{DL_{n}},$ (1)
where $\mathcal{L}_{DL_{n}}$ denotes the total distillation loss of $\mathcal{P}_{n}$ and $\lambda_{n}>0$ is a hyperparameter that controls the relative importance of the cross entropy and distillation losses. Note that the term DL in the above formulation is used generically (without restricting to any specific loss function) and it intuitively captures the differences between the local model of a participant and the models of all its collaborators. The total DL of $\mathcal{P}_{n}$ can in turn computed by aggregating the pairwise distillation losses as follows:
$\mathcal{L}_{DL_{n}}=\sum_{k\in\mathcal{N}\backslash\{n\}}\lambda_{(n,k)}\mathcal{L}_{DL_{(n,k)}},$ (2)
@@
where $\mathcal{L}_{DL_{(n,k)}}$ is the pairwise DL between $\mathcal{P}_{n}$ and $\mathcal{P}_{k}$ and $\lambda_{(n,k)}$ are the weights assigned to the pairwise losses between different participants.
@@
Private Decentralized Learning (PDL): In the PDL framework, the total DL is computed in a privacy-preserving way without leaking the local data of participants. For example, CaPC learning [7] leverages secure multi-party computation, homomorphic encryption, and differential privacy (DP) to securely estimate $\mathcal{L}_{DL_{n}}$. However, CaPC learning requires a semi-trusted third-party privacy guardian to privately aggregate local predictions and the pairwise distillation losses are not accessible. In contrast, CaPriDe learning [25] utilizes fully homomorphic encryption (FHE) to securely compute an approximation of the pairwise DL losses. However, all existing PDL algorithms lack a mechanism to assess the contributions of different participants within the CL framework. Hence, $\lambda_{(n,k)}$ is usually set to $\frac{1}{(N-1)}$ and $\lambda_{n}=\lambda_{0},\leavevmode\nobreak\ \forall\leavevmode\nobreak\ n\in\mathcal{N}$. $\lambda_{0}$ is typically determined through hyperparameter search. We refer to the above scenario as vanilla PDL (VPDL).
@@ 3 Towards a Better Fairness Metric
In [18, 27], collaborative fairness (CF) is defined as follows:
Definition 1.
@@
In a federated system, a high-contribution participant should be rewarded with a better performing local model than a low-contribution participant. Mathematically, fairness can be quantified by the correlation coefficient between the contributions of participants and their respective final model accuracies. [18]
@@
In the above CF Definition 1, the contribution of a participant is typically evaluated based on its standalone local accuracy ($\mathcal{B}_{n}$). Thus, the CF metric attempts to ensure that participants with higher standalone accuracy receive a model with higher final accuracy ($\mathcal{A}_{n}$) compared to that received by other participants. However, the problem with this approach can be illustrated using a simple example. Consider two participants $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ with standalone accuracy of $60\%$ and $80\%$, respectively. Suppose that after collaboration, $\mathcal{P}_{1}$ ($\mathcal{P}_{2}$) receives a model with an accuracy of $70\%$ ($70.1\%$). In this scenario, the system is considered perfectly fair (with a fairness measure of $1$) according to the above CF definition, because $\mathcal{P}_{2}$ receives a model with higher accuracy compared to $\mathcal{P}_{1}$. However, it is obvious that only $\mathcal{P}_{1}$ benefits from this collaboration and achieves a CG of $\mathcal{G}_{1}=10\%$, whereas $\mathcal{P}_{2}$ gains nothing. Strictly speaking, $\mathcal{P}_{1}$ has a negative CG of $\mathcal{G}_{2}=-9.9\%$, but it can always dump the collaboratively trained model and use its own standalone model to achieve zero CG. This scenario cannot be considered as fair, because the participant with the lower contribution benefits significantly without helping the participant with higher contribution.
@@
The problem with CF quantification in Definition 1 is that it overlooks the scenario where the collaboration gain is negative. This is usually not the case in FL with iid settings, because all the participants have similar standalone accuracy and the accuracy of the collaboratively learned model is much higher than any of the individual standalone accuracies. However, in extreme non-iid settings (which is where collaborative fairness is most needed!), the problem of negative CG can be encountered often. Hence, there is a need for a better metric to quantify collaborative fairness.
@@
One possible solution is to measure the rewards in terms of CG, instead of the final model accuracies. Consequently, fairness can be quantified by the correlation coefficient between the contributions of participants ($\mathcal{B}_{n}$) and their respective collaboration gains ($\mathcal{G}_{n}$). However, the drawback of this approach is that it is inherently much harder to improve the performance of the higher contribution participant who starts with a much higher standalone accuracy. Going back to the earlier example, if $\mathcal{P}_{1}$ improves from $60\%$ to $70\%$, $\mathcal{P}_{2}$ has to improve its accuracy from $80\%$ to more than $90\%$ to achieve a fairness score of $1$. While it is much easier to satisfy the first condition, it is hard to meet the second condition using any CL algorithm. Hence, the fairness metric based on the correlation between standalone accuracy and CG is likely to be negative for most cases, indicating poor fairness.
@@
A better alternative is to minimize the probability that the CG is negative or zero, i.e., $P(\mathcal{G}_{n}\leq 0)$ should be small. Intuitively, this is possible only when the expected value ($\mu$) of $\mathcal{G}_{n}$ is a large positive value and if the standard deviation ($\nu$) of $\mathcal{G}_{n}$ is small. Mathematically, using the well-known Chebyshev single-tail inequality, we can show that $P(\mathcal{G}_{n}\leq 0)$ is bounded by $\frac{\nu^{2}}{\nu^{2}+\mu^{2}}=\frac{1}{(1+(\mu/\nu)^{2})}$, provided $\mu>0$. Thus, minimizing $P(\mathcal{G}_{n}\leq 0)$ requires maximizing $\mu$ and minimizing $\nu$. Let the mean collaboration gain (MCG) and collaboration gain spread (CGS) be defined as the sample mean and standard deviation of CG across participants, i.e.,
@@
Hence, an ideal decentralized learning scheme should maximize MCG (achieve better utility), while at the same time minimize CGS (ensure that all participants benefit fairly). This minimizes the likelihood of a participant being negatively impacted by the collaboration (having negative CG).

### v2

@@ Abstract
Collaborative learning (CL) enables multiple participants to jointly train machine learning (ML) models on decentralized data sources without raw data sharing. While the primary goal of CL is to maximize the expected accuracy gain for each participant, it is also important to ensure that the gains are fairly distributed: no client should be negatively impacted, and gains should reflect contributions. Most existing CL methods require central coordination and focus only on gain maximization, overlooking fairness. In this work, we first show that the existing measure of collaborative fairness based on the correlation between accuracy values without and with collaboration has drawbacks because it does not account for negative collaboration gain. We argue that maximizing mean collaboration gain (MCG) while simultaneously minimizing the collaboration gain spread (CGS) is a fairer alternative. Next, we propose the CYCle protocol that enables individual participants in a private decentralized learning (PDL) framework to achieve this objective through a novel reputation scoring method based on gradient alignment between the local cross-entropy and distillation losses. We further extend the CYCle protocol to operate on top of gossip-based decentralized algorithms such as Gossip-SGD. We also theoretically show that CYCle performs better than standard FedAvg in a two-client mean estimation setting under high heterogeneity. Empirical experiments demonstrate the effectiveness of the CYCle protocol to ensure positive and fair collaboration gain for all participants, even in cases where the data distributions of participants are highly skewed. The code can be found at https://github.com/tnurbek/cycle.
@@ 1 Introduction
Collaborative learning (CL) refers to a framework where several entities can work together by pooling their resources to achieve a common machine learning (ML) objective without sharing raw data. This approach offers particular advantages in domains like healthcare and finance that require access to extensive datasets, which can be difficult to acquire for any single entity due to the costs involved or privacy regulations such as HIPAA (Centers for Medicare & Medicaid Services, 1996) and GDPR (European Parliament & Council of the European Union, 2016; Albrecht, 2016). Federated learning (FL) (McMahan et al., 2017) is a specific case of CL that enables multiple entities to collectively train a shared model by sharing their respective model parameters or gradients with a central server for aggregation. However, FL methods can lead to information leakage associated with sharing gradients/parameter updates with an untrusted third-party server that can potentially carry out reconstruction attacks (Zhu et al., 2019; Zhao et al., 2020). Furthermore, most CL algorithms assume equal contributions from all participants. When this assumption is violated, there is little incentive for collaboration because the gains are not distributed fairly. To avoid these pitfalls, a fully decentralized learning algorithm that fairly rewards the contributions of individual participants is required.
Recent works have attempted to tackle the challenge of collaborative fairness (CF) in FL settings. Shapley value (SV) (Shapley et al., 1953) is a common choice for estimating participants’ marginal utility, but its computation and communication costs are high. To mitigate this problem, variants of SV have been proposed in (Wang et al., 2020; Kumar et al., 2022; Xu et al., 2021; Tastan et al., 2024). But these methods only focus on FL with a central server. Existing private decentralized learning (PDL) algorithms such as CaPriDe learning (Tastan & Nandakumar, 2023), CaPC (Choquette-Choo et al., 2021), and Cronus (Chang et al., 2019) emphasize confidentiality, privacy, and utility, while ignoring collaborative fairness.
In this work, we aim to bridge this significant gap by designing the CYCle protocol for knowledge sharing in the PDL setting. We focus on three goals: maximizing the mean accuracy gain across all participants, ensuring that no participant suffers performance degradation due to collaboration, and that the accuracy gains are evenly distributed. To achieve these goals, we make the following contributions:
  • We analyze the collaborative fairness metric based on the correlation coefficient between the contributions of participants and their respective final accuracies, and show that it fails in cases where the collaboration gain is negative.
@@
  • We extend the CYCle protocol to operate seamlessly on top of gossip-based decentralized algorithms such as Gossip-SGD, enabling fair collaboration.
@@ 2 Related Work and Background
Several challenges associated with CL have been addressed recently, including privacy (Dwork et al., 2014; Choquette-Choo et al., 2021; Chang et al., 2019; Tastan & Nandakumar, 2023; Kairouz et al., 2021; Erlingsson et al., 2020; McMahan et al., 2018), confidentiality (Tastan & Nandakumar, 2023; Choquette-Choo et al., 2021; Chang et al., 2019), communication efficiency (McMahan et al., 2017), robustness (Lakshminarayanan et al., 2017; Athalye et al., 2018; Bagdasaryan et al., 2020), and fairness (Xu et al., 2021; Zhou et al., 2021; Lyu et al., 2020). Most existing works focus only on FL algorithms with central orchestration such as FedAvg (McMahan et al., 2017). Due to the potential harm caused by gradient leakage in FL (Zhao et al., 2020; Zhu et al., 2019), which can result in the disclosure of sensitive user data, there is a growing interest in decentralized learning algorithms that do not require centralized orchestration. Focus has been mainly on ensuring confidentiality, privacy, and utility in decentralized learning (Tastan & Nandakumar, 2023; Choquette-Choo et al., 2021; Chang et al., 2019). While frameworks like Cronus (Chang et al., 2019) offer robustness and privacy in decentralized settings, they do not address fairness, which is central to our work. To our knowledge, CYCle is the first to explicitly target collaborative fairness in decentralized learning.
To achieve collaborative fairness in CL, it is critical to evaluate marginal contributions of participants (Jia et al., 2019; Xu et al., 2021; Shi et al., 2022; Jiang et al., 2023; Tastan et al., 2024; Donahue & Kleinberg, 2021). Because data distributions are often heterogeneous, some clients may contribute more valuable data, yet vanilla FL gives all participants the same global model, discouraging collaboration and hindering adoption. While Shapley Value (SV) can be used for data valuation, SV computation is expensive and hard to employ (Shapley et al., 1953). In CGSV (Xu et al., 2021), cosine similarity between local and global parameter updates is used to approximate the SV. The same cosine similarity approach is used in RFFL (Xu & Lyu, 2021) for reputation scoring. Kumar et al. (2022) used logistic regression models as proxies for client data and utilized an ensemble of them to approximate the SVs.
Compared to prior work that either assumes centralized orchestration or overlooks fairness, CYCle introduces a fully decentralized protocol that ensures fair collaboration by regulating knowledge transfer based on participant reputations. It departs from correlation-based fairness metrics that fail under negative collaboration gain and instead proposes a metric specifically designed for fair collaborative learning. CYCle extends naturally to Gossip-SGD, and we show both empirically and theoretically that it achieves better alignment between contribution and reward, particularly under data heterogeneity.
@@ 2.1 Preliminaries
Notations.
Let $\tilde{{\mathcal{M}}}_{\theta}:{\mathcal{X}}\rightarrow\mathcal{Y}$ be a supervised classifier parameterized by $\theta$, where $\mathcal{X}\subseteq\mathbb{R}^{d}$ and $\mathcal{Y}=\{1,2,\cdots,M\}$ denote the input and label spaces, respectively, $d$ is the input dimensionality, and $M$ is the number of classes. Let $\mathcal{M}_{\theta}:\mathcal{X}\rightarrow\mathcal{Z}$ denote a mapping from the input to the logits space ($\mathcal{Z}\subset\mathbb{R}^{M}$) and $\sigma_{T}$ be a softmax function (with temperature parameter $T$) that maps the logits into a probability distribution ${\bm{p}}=\sigma_{T}({\bm{z}})$ over the $M$ classes. The sample is eventually assigned to the class with the highest probability. Given an input sample ${\bm{x}}\in\mathcal{X}$ and its ground truth label $y\in\mathcal{Y}$, let $\mathcal{L}_{CE}$ denote the cross-entropy (CE) loss based on the prediction $\sigma_{T}(\mathcal{M}_{\theta}({\bm{x}}))$ and $y$. We denote ${\mathcal{L}}_{DL_{(i,j)}}$ as the pairwise distillation loss between two models $\mathcal{M}_{\theta_{i}}$ and $\mathcal{M}_{\theta_{j}}$. For example, the distillation loss (DL) could be computed as KL divergence between $\sigma_{T}(\mathcal{M}_{\theta_{i}}({\bm{x}}))$ and $\sigma_{T}(\mathcal{M}_{\theta_{j}}({\bm{x}}))$. Note that other types of distillation losses (Gou et al., 2021) can also be used in lieu of KL divergence.
Decentralized Learning.
Let $N$ be the number of collaborating participants and $\mathcal{N}=\{1,2,...,N\}$. We assume that each participant $\mathcal{P}_{n},n\in\mathcal{N}$ has its own local training dataset $\mathcal{D}_{n}$ and a hold-out validation set $\tilde{\mathcal{D}}_{n}$. Based on the local training set, each participant can learn a local ML model $\mathcal{M}_{\theta_{n}}$ by minimizing the local cross-entropy loss $\mathcal{L}_{CE_{n}}$. The validation accuracy of this standalone local model $\mathcal{M}_{\theta_{n}}$ on $\tilde{\mathcal{D}}_{n}$ is denoted as $\mathcal{B}_{n}$. The goal in a typical FL algorithm is to build a single global model $\mathcal{M}_{\theta_{*}}$ that has a better accuracy than the standalone accuracy of all the local models. In contrast, the goal of each participant in decentralized learning is to obtain a better local model $\mathcal{M}_{\theta_{n}^{*}}$, which has a higher local validation accuracy $\mathcal{A}_{n}$ after collaboration. In other words, each participant $\mathcal{P}_{n}$ aims to maximize its collaborative gain (CG), which is defined as $\mathcal{G}_{n}=(\mathcal{A}_{n}-\mathcal{B}_{n})$. This is typically achieved by minimizing the following objective:
$\mathcal{L}_{n}=\mathcal{L}_{CE_{n}}+\lambda_{n}\mathcal{L}_{DL_{n}},$ (1)
where $\mathcal{L}_{DL_{n}}$ denotes the total distillation loss of $\mathcal{P}_{n}$ and $\lambda_{n}>0$ is a hyperparameter that controls the relative importance of the cross-entropy and distillation losses. Note that the term DL in the above formulation is used generically (without restricting to any specific loss function), and it intuitively captures the differences between the local model of a participant and the models of all its collaborators. The total DL of $\mathcal{P}_{n}$ can in turn be computed by aggregating the pairwise distillation losses as follows:
@@
where $\mathcal{L}_{DL_{(n,k)}}$ is the pairwise DL between $\mathcal{P}_{n}$ and $\mathcal{P}_{k}$, and $\lambda_{(n,k)}$ are the weights assigned to the pairwise losses between different participants.
Private Decentralized Learning (PDL).
@@
In the PDL framework, the total DL is computed in a privacy-preserving way without leaking the local data of participants. For example, CaPC learning (Choquette-Choo et al., 2021) leverages secure multi-party computation, homomorphic encryption, and differential privacy (DP) to securely estimate $\mathcal{L}_{DL_{n}}$. However, CaPC learning requires a semi-trusted third-party privacy guardian to privately aggregate local predictions, and the pairwise distillation losses are not accessible. In contrast, CaPriDe learning (Tastan & Nandakumar, 2023) utilizes fully homomorphic encryption (FHE) to securely compute an approximation of the pairwise DL losses. However, all existing PDL algorithms lack a mechanism to assess the contributions of different participants within the CL framework. Hence, $\lambda_{(n,k)}$ is usually set to $\frac{1}{(N-1)}$ and $\lambda_{n}=\lambda_{0},~\forall~n\in\mathcal{N}$. $\lambda_{0}$ is typically determined through hyperparameter search. We refer to the above scenario as vanilla PDL (VPDL).
Figure 2: Illustration of confidential and private decentralized (CaPriDe) learning framework for NN participants, showing the learning process for only participant $\mathcal{P}_{1}$. Here, $\mathcal{D}_{1}=(\mathcal{X}_{1},\mathcal{Y}_{1})=\{{\bm{x}}_{j,1},y_{j,1}\}_{j=1}^{|{\mathcal{D}}_{1}|}$ is the local data of ${\mathcal{P}}_{1}$, $\color[rgb]{0,0,1}\mathcal{E}(\mathcal{X}_{1})$ denotes the collection of encrypted unlabeled samples of $\mathcal{P}_{1}$, and ${\bm{p}}_{(1,l)}$ and ${\bm{z}}_{(1,l)}$ are the prediction probabilities and logits obtained by applying model $\mathcal{M}_{\theta_{l}}$ of participant $\mathcal{P}_{l}$ to ${\mathcal{X}}_{1}$. Black dashed line represents exchange of encrypted data between participants in each round, and blue dashed line denotes a single transfer at the beginning of the protocol. $\bm{d}$} is a decryption method. $\mathcal{L}_{CE}$ and $\mathcal{L}_{DL}$ represent the cross-entropy and distillation losses, respectively. In the context of our paper, $\mathcal{L}_{{DL}_{(1,k)}}\coloneq\widehat{\mathcal{L}}_{{KL}_{1}}-{\bm{d}}\left(\mathcal{E}\left(\widetilde{\mathcal{L}}_{{KL}_{(1,k)}}\right)\right)$.
CaPriDe Learning.
CaPriDe learning framework (Tastan & Nandakumar, 2023), as described above, uses a knowledge distillation approach and homomorphic encryption to enable secure and private computations. Figure 2 illustrates how CaPriDe learning works. Let $\mathcal{E}(\mathcal{X}_{n})=\{\mathcal{E}({\bm{x}}_{j,n})\}_{j=1}^{|{\mathcal{D}}_{n}|}$ be the collection of encrypted input samples of participant $\mathcal{P}_{n}$. Here, the encryption is based on an FHE scheme with the public key of $\mathcal{P}_{n}$. In CaPriDe learning, each party initiates the collaboration by publishing $\mathcal{E}(\mathcal{X}_{n})$ to the other participants. Let ${\bm{z}}_{j,(n,k)}$ and ${\bm{p}}_{j,(n,k)}$ denote the logits vector and probability vector obtained when model $\mathcal{M}_{\theta_{k}}$ of $\mathcal{P}_{k}$ is applied on the input sample ${\bm{x}}_{j,n}$, i.e., the $j^{th}$ training sample of $\mathcal{P}_{n}$. Assume that at round $(t+1),0\leq t<\tau$ of the collaboration, the current model of $\mathcal{P}_{n}$ is $\mathcal{M}_{\theta_{n}^{t}}$. Then, $\mathcal{P}_{n}$ performs one forward pass on its own training set ${\mathcal{D}}_{n}$ to obtain the predictions $\{{\bm{p}}_{j,(n,n)}\}_{j=1}^{|{\mathcal{D}}_{n}|}$. To enable knowledge transfer between participants, CaPriDe learning makes use of the knowledge distillation (KD) approach (Hinton et al., 2015), where a student model learns to mimic the predictions of a teacher model. In CaPriDe learning, there is no designated teacher model, and mutual KD between multiple peer models is used for knowledge transfer. Let $\mathcal{L}_{DL_{(n,k)}}$ denote the pairwise distillation loss between predictions of $\mathcal{P}_{n}$ and $\mathcal{P}_{k}$, which is defined as:
$\mathcal{L}_{DL_{(n,k)}}=\sum_{j=1}^{|{\mathcal{D}}_{n}|}D_{DL}\left({\bm{p}}_{j,(n,n)},{\bm{p}}_{j,(n,k)}\right)=\widehat{{\mathcal{L}}}_{KL_{n}}-{\bm{d}}\left({\mathcal{E}}\left(\widetilde{{\mathcal{L}}}_{KL_{(n,k)}}\right)\right)$ (3)
where $D_{DL}({\bm{p}}_{1},{\bm{p}}_{2})$ denotes the mimic distance between two predictions ${\bm{p}}_{1}$ and ${\bm{p}}_{2}$ and
$\widehat{{\mathcal{L}}}_{KL_{n}}=\sum_{j=1}^{|{\mathcal{D}}_{n}|}\left({\bm{p}}_{j,(n,n)}\cdot\log{\bm{p}}_{j,(n,n)}\right),\qquad{\mathcal{E}}\left(\widetilde{{\mathcal{L}}}_{KL_{(n,k)}}\right)=\sum_{j=1}^{|{\mathcal{D}}_{n}|}{\mathcal{E}}\left({\bm{p}}_{j,(n,n)}\right)\cdot\frac{{\mathcal{E}}\left({\bm{z}}_{j,(n,k)}\right)}{T},$ (4)
where $T$ is a temperature parameter and $\widetilde{{\mathcal{L}}}_{KL_{(n,k)}}$ is encryption-friendly, approximate KL divergence as per CaPriDe learning (Tastan & Nandakumar, 2023).
@@ 3 Towards a Better Fairness Metric
In (Lyu et al., 2020; Xu & Lyu, 2021), collaborative fairness (CF) is defined as follows:
Definition 1 (Collaborative Fairness).
@@
In a federated system, a high-contribution participant should be rewarded with a better performing local model than a low-contribution participant. Mathematically, fairness can be quantified by the correlation coefficient between the contributions of the participants and their respective final model accuracies.
@@
In the above CF Definition 1, the contribution of a participant is typically evaluated based on its standalone local accuracy ($\mathcal{B}_{n}$). Thus, the CF metric attempts to ensure that participants with higher standalone accuracy receive a model with higher final accuracy ($\mathcal{A}_{n}$) compared to that received by other participants. However, the problem with this approach can be illustrated using a simple example. Consider two participants $\mathcal{P}_{1}$ and $\mathcal{P}_{2}$ with standalone accuracy of $60\%$ and $80\%$, respectively. Suppose that after collaboration, $\mathcal{P}_{1}$ ($\mathcal{P}_{2}$) receives a model with an accuracy of $70\%$ ($70.1\%$). In this scenario, the system is considered perfectly fair (with a fairness measure of $1$) according to the above CF definition, because $\mathcal{P}_{2}$ receives a model with higher accuracy compared to $\mathcal{P}_{1}$. However, it is obvious that only $\mathcal{P}_{1}$ benefits from this collaboration and achieves a CG of $\mathcal{G}_{1}=10\%$, whereas $\mathcal{P}_{2}$ gains nothing. Strictly speaking, $\mathcal{P}_{1}$ has a negative CG of $\mathcal{G}_{2}=-9.9\%$, but it can always dump the collaboratively trained model and use its own standalone model to achieve zero CG. This scenario cannot be considered fair, because the participant with the lower contribution benefits significantly without helping the participant with a higher contribution.
@@
The problem with CF quantification in Definition 1 is that it overlooks the scenario where the collaboration gain is negative. This is usually not the case in FL with iid settings, because all the participants have similar standalone accuracy, and the accuracy of the collaboratively learned model is much higher than any of the individual standalone accuracies. However, in extreme non-iid settings (which is where collaborative fairness is most needed!), the problem of negative CG can be encountered often. Hence, there is a need for a better metric to quantify collaborative fairness.
@@
One possible solution is to measure the rewards in terms of CG rather than final model accuracy. Fairness is then quantified by the correlation between participant contributions ($\mathcal{B}_{n}$) and their CGs ($\mathcal{G}_{n}$). However, this metric is problematic: participants with high standalone accuracy are harder to improve. Going back to the earlier example, if $\mathcal{P}_{1}$ improves from $60\%$ to $70\%$, $\mathcal{P}_{2}$ has to improve its accuracy from $80\%$ to more than $90\%$ to achieve a fairness score of $1$. While it is much easier to satisfy the first condition, it is hard to meet the second condition using any CL algorithm. Hence, the fairness metric based on the correlation between standalone accuracy and CG is likely to be negative for most cases, indicating poor fairness.
@@
A better alternative is to minimize the probability that the CG is negative or zero, i.e., $P({\mathcal{G}}_{n}\leq 0)$ should be small. Let $\mu=\mathbb{E}[\mathcal{G}_{n}]$ and $\nu=\sqrt{\mathrm{Var}[\mathcal{G}_{n}]}$ denote the mean and standard deviation of the collaboration gains, respectively. Intuitively, this is possible only when $\mu$ is a large positive value, and if $\nu$ is small. Mathematically, using the well-known Chebyshev single-tail inequality, we can show that $P(\mathcal{G}_{n}\leq 0)$ is bounded by $\frac{\nu^{2}}{\nu^{2}+\mu^{2}}=\frac{1}{(1+(\mu/\nu)^{2})}$, provided $\mu>0$. Thus, minimizing $P(\mathcal{G}_{n}\leq 0)$ requires maximizing $\mu$ and minimizing $\nu$. Let the mean collaboration gain (MCG) and the collaboration gain spread (CGS) be defined as the sample mean and standard deviation of CG across participants, i.e.,
@@
Hence, an ideal decentralized learning scheme should maximize MCG (achieve better utility), while at the same time minimize CGS (ensure that all participants benefit fairly). This minimizes the likelihood of a participant being negatively impacted by the collaboration (having a negative CG).